# Ab initio study of the radii of oxygen isotopes

### v1

@@ II Formalism
[…]
@@
×$\quad\times\frac{\displaystyle\sum_{\vec{c}}\left[\langle\Psi|\rho_{M}(\vec{c}\,)|\Psi\rangle\sum_{i<j}\delta_{\tau_{i}\alpha}\delta_{\tau_{j}\beta}(\bm{n}_{i}-\bm{n}_{j})^{2}\right]}{\displaystyle\sum_{\vec{c}}\langle\Psi|\rho_{M}(\vec{c}\,)|\Psi\rangle}.$ (12)
@@
Since $\rho_{\sigma\tau}(\bm{n})$ is a rank-one operator, that is, $:\rho_{\sigma\tau}^{2}(\bm{n}):=0$, the rank-one operator method Ma:2023ahg allows us to express $\rho_{M}(\vec{c}\,)$ as
@@
×$\times M(s_{L_{t}/2})\cdots M(s_{1})|\Psi_{0}\rangle|.$ (14)
@@ III Results and discussion
In the partial pinhole algorithm, we take $M=4$ for $\rho_{M}$ in Eq.(10). We have verified that the results are largely insensitive to the specific choice of $M$ SM . To enhance computational efficiency, we fix half of the isospin indices in $\rho_{M}$ as protons and the other half as neutrons. This setup ensures that each configuration $\vec{c}$ contributes simultaneously to all three correlation functions $G_{pp}$, $G_{pn}$, and $G_{nn}$.
@@
Table 1: The charge and matter radii (units are fm) calculated by the partial pinhole algorithm using the N3LO interaction, in comparison with available experiment data. Exp. $(e,e)$ Angeli:2013epw : charge radii deduced from electron scattering data. Exp. $(e,e)$, $(p,p)$ Lapoux:2016exf : matter radii deduced from electron scattering and proton scattering data. Exp. $\sigma_{I}$ Ozawa:2000gx : matter radii deduced from interaction cross sections. Exp. $\sigma_{cc}$ Kaur:2022yoh : matter radii deduced from charge changing cross sections.
@@
The extrapolated charge radii are summarized in Table 1, where results using nine additional sets of three-nucleon forces are included to estimate theoretical uncertainties Elhatisari:pre . The results show that 16O and 17O have very similar charge radii, while 18O exhibits a noticeable increase, consistent with experimental data Angeli:2013epw . We note that, although 16O is commonly described as a spherical doubly magic nucleus within shell-model-based approaches, NLEFT simulations predict a distinct tetrahedral arrangement of alpha clusters Epelbaum:2013paa . Supporting this picture, previous NLEFT studies Shen:2022bak demonstrated that the ground state of 12C robustly converges to an equilateral triangular arrangement of alpha clusters independent of the choice of the initial state. By analogy, assuming that 16O similarly forms a stable tetrahedral cluster configuration as indicated by NLEFT calculations, this intrinsic clustering could explain the success of NLEFT in reproducing oxygen isotope charge radii.
@@
In Table 1, we also compare our results with experimental matter radii extracted from electron and proton scattering [$(e,e)$ and $(p,p)$Lapoux:2016exf , interaction cross sections ($\sigma_{I}$zawa:2000gx , and charge-exchange cross sections ($\sigma_{cc}$aur:2022yoh . Unlike charge radii, the extraction of matter radii often involves some degree of model dependence, leading to discrepancies among different experimental methods. For example, matter radii derived from low-energy proton elastic scattering tend to be systematically larger than those extracted from $\sigma_{I}$ and $\sigma_{cc}$ measurements. The most notable case is 20O, where differences up to 0.2 fm are observed among the three different measurements. As discussed in Ref. Lapoux:2016exf , the $(e,e)$ and $(p,p)$ data provide a more reliable descriptions, since the correlations in the target are usually not included in the other methods. From Table 1, we can find our calculated matter radii agree well with the values extracted from $(e,e)$ and $(p,p)$ data, and thus support the statement given in Ref. Lapoux:2016exf .
@@ Acknowledgements.
We are grateful for discussions with Fabian Hildenbrand, Timo Lähde, Dean Lee, Bing-Nan Lu, Yuan-Zhuo Ma, Xiang-Xiang Sun, and Shuang Zhang. This work was supported in part by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101018170), and by the CAS President’s International Fellowship Initiative (PIFI) (Grant No. 2025PD0022). The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC).

### v2

@@ II Formalism
$\exp\left(-\frac{a_{t}C}{2}\rho^{2}\right)=\sqrt{\frac{1}{2\pi}}\int ds~\exp\left(-\frac{s^{2}}{2}+\sqrt{-a_{t}C}s\rho\right),$ (2)
@@
$\quad\times\frac{\displaystyle\sum_{\vec{c}}\left[\langle\Psi|\rho_{M}(\vec{c}\,)|\Psi\rangle\sum_{i<j}\delta_{\tau_{i}\alpha}\delta_{\tau_{j}\beta}(\bm{n}_{i}-\bm{n}_{j})^{2}\right]}{\displaystyle\sum_{\vec{c}}\langle\Psi|\rho_{M}(\vec{c}\,)|\Psi\rangle}.$ (12)
@@
Since $\rho_{\sigma\tau}(\bm{n})$ is a rank-one operator, that is, $\rho_{\sigma\tau}^{2}(\bm{n}):=0$, the rank-one operator method Ma:2023ahg allows us to express $\rho_{M}(\vec{c}\,)$ as
@@
$\times M(s_{L_{t}/2})\cdots M(s_{1})|\Psi_{0}\rangle|.$ (14)
@@ III Results and discussion
In the partial pinhole algorithm, although one could in principle take $M=2$ for $\rho_{M}$ in Eq. (10) as the correlation functions $G_{\alpha\beta}$ in Eq. (7) are two-body observables, such a choice suffers from low computational efficiency. To improve sampling efficiency, we aim for each configuration $\vec{c}$ to contribute simultaneously to all three types of correlations: $G_{pp}$, $G_{pn}$, and $G_{nn}$. This requirement implies that the isospin indices in $\rho_{M}$ must include at least two protons and two neutrons. We therefore adopt $M=4$ and fix half of the indices to be protons and the other half to be neutrons in our calculations. We have further verified that the final results are largely insensitive to the specific choice of $M$, having tested $M=4$, 6, 8, and 10 SM .
@@
Table 1: The charge and matter radii (units are fm) calculated by the partial pinhole algorithm using the N3LO interaction, where the theoretical uncertainties are estimated using ten sets of three-nucleon forces. The available experiment data are also given for comparison. Exp. $(e,e)$ Angeli:2013epw : charge radii deduced from electron scattering data. Exp. $(e,e)$, $(p,p)$ Lapoux:2016exf : matter radii deduced from electron scattering and proton scattering data. Exp. $\sigma_{I}$ Ozawa:2000gx : matter radii deduced from interaction cross sections. Exp. $\sigma_{cc}$ Kaur:2022yoh : matter radii deduced from charge changing cross sections.
@@
The extrapolated charge radii are summarized in Table 1, where results using nine additional sets of three-nucleon forces are included to estimate theoretical uncertainties Elhatisari:pre . Note that the central values and the corresponding errors are shifted compared to the extrapolated ones in Fig. 1. The results show that 16O and 17O have very similar charge radii, while 18O exhibits a noticeable increase, consistent with experimental data Angeli:2013epw . We note that, although 16O is commonly described as a spherical doubly magic nucleus within shell-model-based approaches, NLEFT simulations predict a distinct tetrahedral arrangement of alpha clusters Epelbaum:2013paa . Similar alpha-cluster structures in 8Be Shen:2024qzi , 12Shen:2022bak , and 20Ne Harris:2025zsh have also been predicted by NLEFT in their ground states. Supporting this picture, previous NLEFT studies Shen:2022bak demonstrated that the ground state of 12C robustly converges to an equilateral triangular arrangement of alpha clusters independent of the choice of the initial state. By analogy, assuming that 16O similarly forms a stable tetrahedral cluster configuration as indicated by NLEFT calculations, this intrinsic clustering could explain the success of NLEFT in reproducing oxygen isotope charge radii.
@@
In Table 1, we also compare our results with experimental matter radii extracted from electron and proton scattering [$(e,e)$ and (p,p)(p,p)Lapoux:2016exf , interaction cross sections ($\sigma_{I}$Ozawa:2000gx , and charge-exchange cross sections ($\sigma_{cc}$Kaur:2022yoh . Unlike charge radii, the extraction of matter radii often involves some degree of model dependence, leading to discrepancies among different experimental methods. For example, matter radii derived from low-energy proton elastic scattering tend to be systematically larger than those extracted from $\sigma_{I}$ and $\sigma_{cc}$ measurements. The most notable case is 20O, where differences up to 0.2 fm are observed among the three different measurements. As discussed in Ref. Lapoux:2016exf , the $(e,e)$ and $(p,p)$ data provide a more reliable descriptions, since the correlations in the target are usually not included in the other methods. From Table 1, we can find our calculated matter radii agree well with the values extracted from $(e,e)$ and $(p,p)$ data, and thus support the statement given in Ref. Lapoux:2016exf .
@@ Acknowledgements.
We are grateful for discussions with Fabian Hildenbrand, Timo Lähde, Dean Lee, Bing-Nan Lu, Yuan-Zhuo Ma, Xiang-Xiang Sun, and Shuang Zhang. This work was supported in part by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101018170), and by the CAS President’s International Fellowship Initiative (PIFI) (Grant No. 2025PD0022). The work of SE is supported in part by the Scientific and Technological Research Council of Turkey (TUBITAK project no. 123F464). The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC). Furthermore, the authors gratefully acknowledge the computing time provided on the high-performance computer HoreKa by the National High-Performance Computing Center at KIT (NHR@KIT). This center is jointly supported by the Federal Ministry of Education and Research and the Ministry of Science, Research and the Arts of Baden-Württemberg, as part of the National High-Performance Computing (NHR) joint funding program (https://www.nhr-verein.de/en/our-partners). HoreKa is partly funded by the German Research Foundation (DFG).